OpenStudy (anonymous):

lim of infinity sqrt9x to the 6th power -x all over xsquared +1 I got 3 to be the limit. If thats right, im not sure how I got the answer

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty}\frac{\sqrt{9x^6-1}}{x^2+1}\]

OpenStudy (anonymous):

if so, should be infinity

OpenStudy (anonymous):

no I think the -1 is outside the square root

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty}\frac{\sqrt{9x^6-x}}{x^2+1}\]

OpenStudy (anonymous):

then it simplifies to (3x^3 - 1)/(x^2 + 1)

OpenStudy (anonymous):

Then divide top and bottom by x^3 and you are left with 3/1

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

numerator is a cube, denominator is a square plus how does is simplify?

OpenStudy (anonymous):

nay is right

OpenStudy (anonymous):

aahh yeas you're right

OpenStudy (anonymous):

again you do this with your eyeballs. ignore everything but highest power. the numerator is not a polynomial, but \[\sqrt{x^6}=x^3\] so think "degree 3" denominator is a polynomial of degree 2 bigger numerator means bigger number, so no limit (or it is infinity)

OpenStudy (anonymous):

its square root of 9x to the 6th power - x all over x cubed plus 1

OpenStudy (anonymous):

ooooooooooooooooooh x cubed!!!!

OpenStudy (anonymous):

hmmmm not to give you a hard time... but you wrote squared!

OpenStudy (anonymous):

then you are right good work

OpenStudy (anonymous):

haha yeah I think I solved it with x^3...yeah good job calculus.

OpenStudy (anonymous):

good work to nayeaddo too

OpenStudy (anonymous):

so then a problem like x+x cubed+x to the 5th power all over 1-x squared +x to the fourth power would be infinity???

OpenStudy (anonymous):

lol @ satellite73

OpenStudy (anonymous):

can you maybe write it using the equation tool?

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty}\frac{x+x^3+x^5}{x-x^2+x^4}\] ?

OpenStudy (anonymous):

that is infinity because the degree of the numerator is higher than the denominator

OpenStudy (anonymous):

think about what it would be like if x were 1,000,000 = 10^6

OpenStudy (anonymous):

\[(x + x^3 + x^5)\]/ (1 - x^2 + x^4)

OpenStudy (anonymous):

but solutions is the same regardless.

OpenStudy (anonymous):

or even ten for that matter. suppose you made x = 10 you would have \[\frac{101010}{10101}\]

OpenStudy (anonymous):

numerator is bigger than denominator by a factor of ten. this gets huge as x gets huge